order:1.5g in 100 ml D5W IVPB every 6 hours. Therapeutic range is 100-200 mg/kg/day. patient weighs 110 pounds
Answer #1

Patient weight is 50 kgs (to convert pounds to kgs divide with 2.2)

Therapeutic range is 100-200 mg/kg = 5000-10000 mg per day or 5-10 gms per day

Flow rate = 100\div6 = 16.7 ml per hour

                = 16.7 \div60 = 0.28 or 0.3 ml per minute

If the drop factor of tubing is 20 gtts per 1 ml then the drops per minute = 0.3 \times 20 = 6 drops per minute